# thanks: Corresponding Author

### v1

@@ I Introduction
The widths of the decay channels depend on numerous input parameters of the molecules $\mathcal{M}_{\mathrm{b}}$ and $\mathcal{M}_{\mathrm{c}}$, and of final-state mesons. The masses and couplings of $\mathcal{M}_{\mathrm{b}}$ and $\mathcal{M}_{\mathrm{c}}$ are object of the present studies. The parameters of the conventional mesons are known from experimental measurements or were found using different theoretical methods. Decisive quantities which should be determined are the strong couplings at the, for instance, vertices $\mathcal{M}_{\mathrm{b}}\eta_{b}B_{c}^{-}$, $\mathcal{M}_{\mathrm{c}}\eta_{c}B_{c}^{+}$ and $\mathcal{M}_{\mathrm{c}}J/\psi B_{c}^{\ast+}$. They describe the strong interaction of the molecule with ordinary final-state mesons and can be estimated by means of the QCD three-point sum rule method that allows one to evaluate relevant form factors.
This paper is organized in the following way: In Sec. II, we compute the masses and current couplings of the scalar molecules $\mathcal{M}_{\mathrm{b}}$ and $\mathcal{M}_{\mathrm{c}}$. The width of the molecule $\mathcal{M}_{\mathrm{b}}$ is computed in Sec. III. The full width of the structure $\mathcal{M}_{\mathrm{c}}$ saturated by the aforementioned modes is determined in section IV. We make our conclusions in the last part of the article V.
@@
$\widetilde{J}(x)=\overline{c}_{a}(x)i\gamma_{5}c_{a}(x)\overline{c}_{b}(x)i\gamma_{5}b_{b}(x).$ (2)
@@ II.1 Parameters of the molecule $\mathcal{M}_{\mathrm{b}}$
The spectral density $\rho^{\mathrm{OPE}}(s)$ is found as an imaginary part of the function $\Pi^{\mathrm{OPE}}(p^{2})$. Because in the current paper we consider only the perturbative and dimension-four nonperturbative contributions $\sim\langle\alpha_{s}G^{2}/\pi\rangle$ to $\Pi^{\mathrm{OPE}}(p^{2})$, $\rho^{\mathrm{OPE}}(s)$ contains terms $\rho^{\mathrm{pert.}}(s)$ and $\rho^{\mathrm{Dim4}}(s)$. The nonperturbative function $\Pi(M^{2})$ is calculated directly from the correlator $\Pi^{\mathrm{OPE}}(p)$ and embrace effects of terms which are not included into the spectral density.
To carry out the numerical calculations, we have to fix the parameters in Eqs. (8) and (9). The $b$ and $c$ quarks’ masses and gluon condensate $\langle\alpha_{s}G^{2}/\pi\rangle$ are universal quantities. In the current article, we employ
@@
Quark masses $m_{c}$ and $m_{b}$ are calculated in the $\overline{\mathrm{MS}}$ scheme PDG:2024 . The condensate $\langle\alpha_{s}G^{2}/\pi\rangle$ was estimated in Refs. Shifman:1978bx ; Shifman:1978by from studies of different processes.
The parameters $M^{2}$ and $s_{0}$ depend on a analyzing problem and have to satisfy standard restrictions of SR analyses. In the SR method the pole contribution ($\mathrm{PC}$) should dominate in obtained quantities, therefore, in computations we require fulfilment $\mathrm{PC}\geq 0.5$ . Convergence of $\mathrm{OPE}$ is another condition for reliable SR studies. In our case, the correlation function contains only dimension-$4$ term $\Pi^{\mathrm{Dim4}}(M^{2},s_{0})$. Then, the constraint $|\Pi^{\mathrm{Dim4}}(M^{2},s_{0})|\leq 0.05|\Pi(M^{2},s_{0})|$ is enough to ensure convergence of $\mathrm{OPE}$. Last but not least is stability of final results upon variations of $M^{2}$ and $s_{0}$.
@@
meet all these conditions. In fact, at maximal and minimal $M^{2}$ the pole contribution averaged over $s_{0}$ is $\mathrm{PC}\approx 0.52$ and $\mathrm{PC}\approx 0.65$. At $M^{2}=15~\mathrm{GeV}^{2}$ the nonperturbative contribution constitutes approximately $1.5\%$ of the full result. The $\mathrm{PC}$ as afunction of the Borel parameter is presented in Fig. 1, where all lines overshot the border $\mathrm{PC}=0.5$.
@@
The predictions in Eq. (13) amount to SR results at $M^{2}=16.5~\mathrm{GeV}^{2}$ and $s_{0}=277.5~\mathrm{GeV}^{2}$, where $\mathrm{PC}\approx 0.58$, which guaranties the prevalence of $\mathrm{PC}$ in extracted quantites. The ambiguities in Eq. (13) are formed due to choices of $M^{2}$ and $s_{0}$: Uncertainties connected with errors in quark masses and gluon condensate are neglicible.
@@
Figure 3: Mass $\widetilde{m}$ as a function on the parameters $M^{2}$ (left panel), and $s_{0}$ (right panel). The purple triangle shows point $M^{2}=9~\mathrm{GeV}^{2}$ and $s_{0}=109~\mathrm{GeV}^{2}$.
@@ III Full decay width of $\mathcal{M}_{\mathrm{b}}$
In this section we calculate the full decay width of the hadronic molecule $\mathcal{M}_{\mathrm{b}}$. Information on the mass of $\mathcal{M}_{\mathrm{b}}$ permits one to find its decay channels. The process $\mathcal{M}_{\mathrm{b}}\rightarrow\eta_{b}B_{c}^{-}$ is kinematically allowed decay channel of $\mathcal{M}_{\mathrm{b}}$. In fact, the masses $m_{\eta_{b}}=(9398.7\pm 2.0)~\mathrm{MeV}$ and $m_{B_{c}}=(6274.47\pm 0.27\pm 0.17)~\mathrm{MeV}$ PDG:2024 of the final-state mesons establish the threshold $15673~\mathrm{MeV}$ which is less than $m$. This is the dominant mode of $\mathcal{M}_{\mathrm{b}}$, because all of its valence quarks appear in the final-state particles.
Another decay channels of the molecule $\mathcal{M}_{\mathrm{b}}$ are ones generated by annihilation of $b\overline{b}$ quarks in $\mathcal{M}_{\mathrm{b}}$ to $q\overline{q}$ and $s\overline{s}$ pairs. Then initial $b$ and $\overline{c}$ quarks from $\mathcal{M}_{\mathrm{b}}$ and light quarks form pairs of $BD$ mesons with appropriate quantum numbers and charges. Because in the SR method the vacuum expectation value $\langle\overline{b}b\rangle$ of $b$ quarks is replaced by the gluon condensate $\langle\alpha_{s}G^{2}/\pi\rangle$, these processes are subleading modes of the molecule $\mathcal{M}_{\mathrm{b}}$. Nevertheless, total contribution of such channel to the full decay width of $\mathcal{M}_{\mathrm{b}}$ may be sizeable. Here, we are going to take into account decays to mesons $B^{-}\overline{D}^{0}$, $\overline{B}^{0}D^{-}$, $\overline{B}_{s}^{0}D_{s}^{-}$, $B^{\ast-}\overline{D}^{\ast 0}$, $\overline{B}^{\ast 0}D^{\ast-}$, and $\overline{B}_{s}^{\ast 0}D_{s}^{\ast-}$.
@@ III.1 Process $\mathcal{M}_{\mathrm{b}}\rightarrow\eta_{b}B_{c}^{-}$
Mesons mass ($\mathrm{MeV}$) DC ($\mathrm{MeV}$)
@@
Table 1: Masses and decay constants (DC) of the mesons that appear in decays of the hadronic molecules $\mathcal{M}_{b}$ and $\mathcal{M}_{c}$.
@@
The error above is generated by the ambiguities of the coupling $g$ and the masses of the particles $\mathcal{M}_{\mathrm{b}}$ (upper limit), $m_{\eta_{b}}$ and $m_{B_{c}}$.
@@ III.2 Decays of $\mathcal{M}_{\mathrm{b}}$ triggered by $\overline{b}b$ annihilation
$\Pi_{1}^{\mathrm{OPE}}(p,p^{\prime})=\frac{\langle\overline{b}b\rangle}{3}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\mathrm{Tr}\left[\gamma_{5}S_{b}^{ia}(y-x)\right$.
$\left.\times S_{c}^{aj}(x)\gamma_{5}S_{u}^{ji}(-y)\right].$ (35)
To continue calculations, we utilize the relation
$\langle\overline{b}b\rangle=-\frac{1}{12m_{b}}\langle\frac{\alpha_{s}G^{2}}{\pi}\rangle$ (36)
extracted in Ref. Shifman:1978bx using the sum rule method.
@@
$g_{2}\equiv\mathcal{G}_{2}(-m_{D_{s}}^{2},m^{2})=(1.84\pm 0.32)\ \mathrm{GeV}^{-1},$ (40)
@@
$\Pi_{\mu\nu}^{\mathrm{OPE}}(p,p^{\prime})$ $=$ $\frac{\langle\overline{b}b\rangle}{3}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\mathrm{Tr}\left[\gamma_{\mu}S_{b}^{ia}(y-x)\right$. (47)
$\left.\times S_{c}^{aj}(x)\gamma_{\nu}S_{u}^{ji}(-y)\right].$

### v2

@@ I Introduction
The widths of the decay channels depend on numerous input parameters of the molecules $\mathcal{M}_{\mathrm{b}}$ and $\mathcal{M}_{\mathrm{c}}$, and those of final-state mesons. The masses and couplings of $\mathcal{M}_{\mathrm{b}}$ and $\mathcal{M}_{\mathrm{c}}$ are object of the present studies. The parameters of the conventional mesons are known from experimental measurements or were found using different theoretical methods. Decisive quantities which should be determined are the strong couplings at, for instance, the vertices $\mathcal{M}_{\mathrm{b}}\eta_{b}B_{c}^{-}$, $\mathcal{M}_{\mathrm{c}}\eta_{c}B_{c}^{+}$ and $\mathcal{M}_{\mathrm{c}}J/\psi B_{c}^{\ast+}$. They describe the strong interaction of the molecule with ordinary final-state mesons and can be estimated by means of the QCD three-point sum rule method that allows one to evaluate relevant form factors.
This paper is organized in the following way: In Sec. II, we compute the masses and current couplings of the scalar molecules $\mathcal{M}_{\mathrm{b}}$ and $\mathcal{M}_{\mathrm{c}}$. The width of the molecule $\mathcal{M}_{\mathrm{b}}$ is calculated in Sec. III. The full width of the structure $\mathcal{M}_{\mathrm{c}}$ saturated by the aforementioned modes is determined in section IV. We make our conclusions in the last part of the article V.
@@
$\widetilde{J}(x)=\overline{c}_{a}(x)i\gamma_{5}c_{a}(x)\overline{b}_{b}(x)i\gamma_{5}c_{b}(x).$ (2)
@@ II.1 Parameters of the molecule $\mathcal{M}_{\mathrm{b}}$
We tacitly assume in Eq. (4) that the physical side of SR can be approximated by a single pole term. But in the multiquark systems two-meson reducible terms also contribute to $\Pi^{\mathrm{Phys}}(p)$ Kondo:2004cr ; Lee:2004xk . The reason is that the current $J(x)$ interacts not only with a molecule $\mathcal{M}_{\mathrm{b}}$, but also with the relevant two-meson continuum. Such interaction, properly included into analysis, leads to a finite width of the hadronic molecule and modifies the quark propagator Agaev:2022ast . It was numerously demonstrated that two-meson contributions can be taken into account by rescaling the current coupling $\Lambda$ but keeping stable the mass $m$ of an exotic four-quark meson of interest. Computations prove that these effects are small and do not overshoot ambiguities of SR analysis: Even for diquark-antidiquark systems with the widths of a few hundred $\mathrm{MeV}$ corresponding modifications amount to additional $5-7\%$ ambiguities in the current couplings Agaev:2018vag ; Sundu:2018nxt , whereas SRs generate uncertainties around of $10\%$ and higher. For the molecules $\mathcal{M}_{\mathrm{c}}$ and $\mathcal{M}_{\mathrm{b}}$ with the widths $70-90~\mathrm{MeV}$ the two-meson contaminations most likely are small, therefore in $\Pi^{\mathrm{Phys}}(p)$ we use the zero-width single-pole approximation.
@@
The spectral density $\rho^{\mathrm{OPE}}(s)$ is found as an imaginary part of the function $\Pi^{\mathrm{OPE}}(p^{2})$. In the current paper we consider perturbative and dimension-four $\sim\langle\alpha_{s}G^{2}/\pi\rangle$ and six $\sim\langle g_{s}^{3}G^{3}\rangle$ contributions to $\Pi^{\mathrm{OPE}}(p^{2})$, therefore $\rho^{\mathrm{OPE}}(s)$ contains terms $\rho^{\mathrm{pert.}}(s)$, $\rho^{\mathrm{Dim4}}(s)$, and $\rho^{\mathrm{Dim6}}(s)$. The nonperturbative function $\Pi(M^{2})$ is calculated directly from the correlator $\Pi^{\mathrm{OPE}}(p)$ and embrace effects of terms which are not included into the spectral density.
To carry out the numerical calculations, we have to fix the parameters in Eqs. (8) and (9). The $b$ and $c$ quarks’ masses and gluon condensates $\langle\alpha_{s}G^{2}/\pi\rangle$ and $\langle g_{s}^{3}G^{3}\rangle$ are universal quantities. In the current article, we employ
@@
$\langle g_{s}^{3}G^{3}\rangle=(0.57\pm 0.29)~\mathrm{GeV}^{6}$ (12)
Quark masses $m_{c}$ and $m_{b}$ are calculated in the $\overline{\mathrm{MS}}$ scheme PDG:2024 . The condensates $\langle\alpha_{s}G^{2}/\pi\rangle$ and $\langle g_{s}^{3}G^{3}\rangle$ were estimated in Refs. Shifman:1978bx ; Shifman:1978by ; Narison:2015nxh from studies of different processes.
Figure 1: Contributions of different terms to $\Pi(M^{2},s_{0})$ as functions of $M^{2}$ at fixed $s_{0}=277.5~\mathrm{GeV}^{2}$.
The parameters $M^{2}$ and $s_{0}$ depend on a analyzing problem and have to satisfy standard restrictions of SR analyses. In the SR method the pole contribution ($\mathrm{PC}$) should dominate in obtained quantities, therefore, in computations we require fulfilment $\mathrm{PC}\geq 0.5$ . Convergence of $\mathrm{OPE}$ is another condition for reliable SR studies. In our case, the correlation function contains dimension-$4$ and -$6$ terms $\Pi^{\mathrm{Dim4}}(M^{2},s_{0})$ and $\Pi^{\mathrm{Dim6}}(M^{2},s_{0})$. Then, the constraint $|\Pi^{\mathrm{Dim4}}(M^{2},s_{0})+\Pi^{\mathrm{Dim6}}(M^{2},s_{0})|\leq 0.05|\Pi(M^{2},s_{0})|$ is enough to ensure convergence of $\mathrm{OPE}$. Last but not least is stability of final results upon variations of $M^{2}$ and $s_{0}$.
@@
meet all these conditions. In fact, at maximal and minimal $M^{2}$ the pole contribution averaged over $s_{0}$ is $\mathrm{PC}\approx 0.52$ and $\mathrm{PC}\approx 0.65$. The nonperturbative contributions are negative and at $M^{2}=15~\mathrm{GeV}^{2}$ $|\Pi^{\mathrm{Dim4}}(M^{2},s_{0})+\Pi^{\mathrm{Dim6}}(M^{2},s_{0})|$ constitutes approximately $2\%$ of the full result. In Fig. 1 we plot the ratio
$R_{\mathrm{pert.(N)}}(M^{2},s_{0})=\left|\frac{\Pi^{\mathrm{pert.(DimN)}}(M^{2},s_{0})}{\Pi(M^{2},s_{0})}\right|,$ (14)
for the perturbative and nonperturbative components of $\Pi(M^{2},s_{0})$. As is seen, contributions of the dimension-$6$ terms are negligibly small, therefore in what follows we truncate $\mathrm{OPE}$ at dimension-$4$ terms. The $\mathrm{PC}$ as a function of the Borel parameter is presented in Fig. 2, where all lines overshot the border $\mathrm{PC}=0.5$.
@@
The predictions in Eq. (15) amount to SR results at $M^{2}=16.5~\mathrm{GeV}^{2}$ and $s_{0}=277.5~\mathrm{GeV}^{2}$, where $\mathrm{PC}\approx 0.58$, which guaranties the prevalence of $\mathrm{PC}$ in extracted quantities. The ambiguities in Eq. (15) are formed due to choices of $M^{2}$ and $s_{0}$: Uncertainties connected with errors in quark masses and gluon condensate are negligible.
@@
Figure 4: Mass $\widetilde{m}$ as a function on the parameters $M^{2}$ (left panel), and $s_{0}$ (right panel). The triangle shows point $M^{2}=9~\mathrm{GeV}^{2}$ and $s_{0}=109~\mathrm{GeV}^{2}$.
@@ III Full decay width of $\mathcal{M}_{\mathrm{b}}$
In this section we calculate the full decay width of the hadronic molecule $\mathcal{M}_{\mathrm{b}}$. Information on the mass of $\mathcal{M}_{\mathrm{b}}$ permits one to find its decay channels. The process $\mathcal{M}_{\mathrm{b}}\rightarrow\eta_{b}B_{c}^{-}$ is kinematically allowed decay channel of $\mathcal{M}_{\mathrm{b}}$. In fact, the masses $m_{\eta_{b}}=(9398.7\pm 2.0)~\mathrm{MeV}$ and $m_{B_{c}}=(6274.47\pm 0.27\pm 0.17)~\mathrm{MeV}$ PDG:2024 of the final-state mesons establish the threshold $15673~\mathrm{MeV}$ which is less than central value of $m=15728~\mathrm{MeV}$. This is the dominant mode of $\mathcal{M}_{\mathrm{b}}$, because all of its valence quarks appear in the final-state particles.
It is interesting that $m=15728~\mathrm{MeV}$ is rather close to $\eta_{b}B_{c}^{-}$ threshold and, due to uncertainties in estimation of $m$, in its lower value $m=15638~\mathrm{MeV}$ lies below it. In other words the scalar molecule $\mathcal{M}_{\mathrm{b}}$ can be considered as a bound state of the mesons $\eta_{b}$ and $B_{c}^{-}$. Then the dominant decay channel $\mathcal{M}_{\mathrm{b}}\rightarrow\eta_{b}B_{c}^{-}$ becomes kinematically forbidden for $\mathcal{M}_{\mathrm{b}}$. But this does not mean that it is stable against strong-interaction decays. In fact, there is an alternative mechanism for its transformation to ordinary particles.
The alternative decay channels of the molecule $\mathcal{M}_{\mathrm{b}}$ are ones generated by annihilation of $b\overline{b}$ quarks in $\mathcal{M}_{\mathrm{b}}$ to $q\overline{q}$ and $s\overline{s}$ pairs. Afterwards, initial $b$ and $\overline{c}$ quarks from $\mathcal{M}_{\mathrm{b}}$ and light quarks form pairs of $B_{(s)}^{(\ast)}D_{(s)}^{(\ast)}$ mesons with appropriate quantum numbers and charges. In the SR method in corresponding correlation functions we relate the vacuum expectation value $\langle\overline{b}b\rangle$ of $b$ quarks to the gluon condensate $\langle\alpha_{s}G^{2}/\pi\rangle$, therefore these processes are subleading modes of the molecule $\mathcal{M}_{\mathrm{b}}$. Nevertheless, total contribution of such channel to the full decay width of $\mathcal{M}_{\mathrm{b}}$ may be sizeable. Here, we are going to take into account decays to mesons $B^{-}\overline{D}^{0}$, $\overline{B}^{0}D^{-}$, $\overline{B}_{s}^{0}D_{s}^{-}$, $B^{\ast-}\overline{D}^{\ast 0}$, $\overline{B}^{\ast 0}D^{\ast-}$, and $\overline{B}_{s}^{\ast 0}D_{s}^{\ast-}$.
@@ III.1 Process $\mathcal{M}_{\mathrm{b}}\rightarrow\eta_{b}B_{c}^{-}$
Mesons mass ($\mathrm{MeV}$) D.C. ($\mathrm{MeV}$)
@@
Table 1: Masses and decay constants (D.C.) of the mesons that appear in decays of the hadronic molecules $\mathcal{M}_{b}$ and $\mathcal{M}_{c}$.
@@
The error above is generated by the ambiguities of the coupling $g$ and the masses of the particles $\mathcal{M}_{\mathrm{b}}$ (upper limit), $\eta_{b}$ and $B_{c}^{-}$.
@@ III.2 Decays of $\mathcal{M}_{\mathrm{b}}$ triggered by $\overline{b}b$ annihilation
$\Pi_{1}^{\mathrm{OPE}}(p,p^{\prime})=\frac{1}{3}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle\overline{b}b\rangle$
$\times\mathrm{Tr}\left[\gamma_{5}S_{b}^{ia}(y-x)S_{c}^{aj}(x)\gamma_{5}S_{u}^{ji}(-y)\right].$ (37)
The correlation function $\Pi_{1}^{\mathrm{OPE}}(p,p^{\prime})$ contains three quark propagators and vacuum condensate $\langle\overline{b}b\rangle$, and differs from standard correlators, for example, from Eq. (24), which depends on four quark propagators. To calculate $\Pi_{1}^{\mathrm{OPE}}(p,p^{\prime})$, as usual, we contract heavy and light quark fields. Because the pair of mesons $B^{-}\overline{D}^{0}$ contains only one $b$-quark field, remaining pair of $\overline{b}b$ quarks in the molecule $\mathcal{M}_{\mathrm{b}}$ form a local vacuum condensate. This is a reason for appearance of the factor $\langle\overline{b}b\rangle$ in Eq. (37). Stated differently, $\langle\overline{b}b\rangle$ emerges as a substituent of a quark propagator and has to be treated on equal footing with them.
To continue calculations, we employ the heavy and light quark propagators and use their standard formulas. We utilize also the relation between the heavy quark and gluon condensates to express $\Pi_{1}^{\mathrm{OPE}}(p,p^{\prime})$ in terms of known parameters. The available expression for $\langle\overline{b}b\rangle$ reads Shifman:1978bx ; Generalis:1983hb ; Bagan:1985zp
$m_{b}\langle\overline{b}b\rangle$ $=$ $-\frac{1}{12}\langle\frac{\alpha_{s}G^{2}}{\pi}\rangle+\frac{1}{m_{b}^{2}}\langle\frac{\alpha_{s}G^{3}}{\pi}\rangle$ (38)
$\times\left(-\frac{1}{48}+\frac{13}{720}\right)+\cdots.$
In our analysis we use only the first term in Eq. (38) extracted in Ref. Shifman:1978bx , because next ones are suppressed by additional powers of $m_{b}^{-1}$ and can be neglected.
@@
$g_{2}\equiv\mathcal{G}_{2}(-m_{D_{s}}^{2},m^{2})=(1.84\pm 0.32)\ \times 10^{-2}\ \mathrm{GeV}^{-1},$ (42)
@@
$\Pi_{\mu\nu}^{\mathrm{OPE}}(p,p^{\prime})=\frac{1}{3}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle\overline{b}b\rangle$
$\times\mathrm{Tr}\left[\gamma_{\mu}S_{b}^{ia}(y-x)S_{c}^{aj}(x)\gamma_{\nu}S_{u}^{ji}(-y)\right].$ (49)
@@
It has been emphasized above that at lower mass ($\mathrm{l.m.}$) $m=15638~\mathrm{MeV}$ the molecule $\mathcal{M}_{\mathrm{b}}$ is stable against the dominant decay channel. As a result, its full decay width is formed due to subleading processes. To evaluate $\Gamma\left[\mathcal{M}_{\mathrm{b}}\right]|_{\mathrm{l.m.}}$, we have repeated calculations of the current subsection with $m=15638~\mathrm{MeV}$. Our prediction for the width of the molecule $\mathcal{M}_{\mathrm{b}}$ reads
$\Gamma\left[\mathcal{M}_{\mathrm{b}}\right]|_{\mathrm{l.m.}}=(49\pm 6)~\mathrm{MeV}.$ (60)